calculate the hydronium and hydroxide concentrations of a solution with a pH of 4.5, one with a pH of 7, and another with a pH of 12.63
Answer #1

As we know that, pH=-log(H3O+)

[H3O+]=10^-pH and [OH-]=10^-pOH

And [H3O+][OH-]=1x10^-14

pH+pOH=14

pH=14-pOH

1. pH=4.5, pOH=14-4.5=14-4.5= 9.5

[H3O+]= 10^-4.5= 3.16x10^-5 M

[OH-]=10^-9.5= 3.16x10^-10 M

2. pH=7

[H3O+]= 10^-7 M

pOH=14-7=7

And [OH-]=10^-7 M

pH=12.63

pOH=14-12.63= 1.37

[H3O+]= 10^-12.63= 2.34x10^-13 M

[OH-]= 10^1.37=4.26x10^-2 M